# Value-oriented forecast reconciliation for renewables in electricity markets

### v1

@@ Abstract
Forecast reconciliation is considered an effective method for achieving coherence and improving forecast accuracy. However, the value of reconciled forecasts in downstream decision-making tasks has been mostly overlooked. In a multi-agent setup with heterogeneous loss functions, this oversight may lead to unfair outcomes, hence resulting in conflicts during the reconciliation process. To address this, we propose a value-oriented forecast reconciliation approach that focuses on the forecast value for individual agents. Fairness is ensured through the use of a Nash bargaining framework. Specifically, we model this problem as a cooperative bargaining game, where each agent aims to optimize their own gain while contributing to the overall reconciliation process. We then present a primal-dual algorithm for parameter estimation based on empirical risk minimization. From an application perspective, we consider an aggregated wind energy trading problem, where profits are distributed using a weighted allocation rule. We demonstrate the effectiveness of our approach through several numerical experiments, showing that it consistently results in increased profits for all agents involved.
@@ 1 Introduction
The geographically distributed nature of renewable energy sources, such as wind and solar, inherently creates a hierarchical structure for power generation. It is acknowledged that forecast accuracy can be improved by leveraging the spatial-temporal dependencies inherent in geographically distributed data (Tastu et al.,, 2013). However, various agents at different levels of aggregation in power systems and electricity markets, for instance, the portfolio and individual power generation assets, often independently generate their own forecasts. This can lead to incoherence, i.e., the forecast directly produced at the portfolio level is not necessarily equal to the sum of the forecasts for individual assets. Given that forecasts are frequently used in decision-making, incoherence across aggregation levels may eventually lead to dilemmas among decision-makers. For that, forecast reconciliation is applied, which is a post-forecasting process that involves transforming independently generated incoherent base forecasts into coherent forecasts that satisfy the constraints imposed by the hierarchical structures (Athanasopoulos et al.,, 2024).
@@
Single-level reconciliation approaches, such as bottom-up, top-down, and middle-out approaches, have been in use for decades, dating back at least to Edwards and Orcutt, (1969) and Zellner, (1969). They require forecasts at a chosen level, which are then combined to generate forecasts for the remaining levels in the hierarchy. Hyndman et al., (2011) introduced a combination approach that leverages forecasts across the entire hierarchy. This was later extended by Van Erven and Cugliari, (2015), Hyndman et al., (2016), and Wickramasuriya et al., (2019) through employing various projection matrices (Panagiotelis et al.,, 2021). In parallel, Ben Taieb and Koo, (2019) developed an end-to-end reconciliation approach, which learns the projection matrices via machine learning. Forecast reconciliation in the energy sector has been employed since Van Erven and Cugliari, (2015), and most often brings the added benefit of enhanced forecast accuracy (Zhang and Dong,, 2018; Jeon et al.,, 2019).
@@
While previous studies have substantially advanced forecast reconciliation through various methods and techniques, nearly all of them focus on leveraging hierarchical structures to improve forecast accuracy. However, it is believed that in addition to evaluating the quality of forecasts, their value in downstream decision-making processes should also be assessed (Murphy,, 1993; Theodorou et al.,, 2025), in terms of profits or costs. Although increasing advancements focus on developing forecasting products that maximize the value in decision-making, particularly within the energy industry (Carriere and Kariniotakis,, 2019; Stratigakos et al.,, 2022; Zhang et al.,, 2024), drawing inspiration from the pioneering research of Bertsimas and Kallus, (2020); Elmachtoub and Grigas, (2022), the value in forecast reconciliation remains largely under-investigated, with only a few notable exceptions. For instance, Sanguri et al., (2022) assessed the effectiveness of coherent forecasts in enhancing the value of port operations and empirically demonstrated that coherent forecasts improved decision-making at ports. In parallel, Stratigakos et al., (2024) proposed a method for combining multiple probabilistic forecasts using a linear opinion pool, with a particular emphasis on minimizing the expected costs associated with subsequent decision-making. However, all of these studies evaluate the value in forecast reconciliation from a centralized perspective, neglecting the fact that the agents involved may have heterogeneous loss (or cost) functions.
@@
When agents have heterogeneous cost functions, the primary challenge lies in achieving consensus on the reconciliation approach before actually reconciling their forecasts. Since the commonly used reconciliation approaches aim to optimize an overall loss function, for instance, the mean squared error function (Van Erven and Cugliari,, 2015; Hyndman et al.,, 2016), some agents involved may even find that reconciled forecasts lead to the increase of their own costs, even though overall costs are reduced after reconciliation. As a result, the agents may perceive the reconciliation process as unfair, leading to disagreements on the reconciliation procedure. In fact, Van Erven and Cugliari, (2015) incorporated weighting factors for forecasts at each level during reconciliation. However, these weights were assigned ad hoc, and the impact of such weighting on downstream decision-making outcomes was not explored. Therefore, how to account for the value of forecasts at each level and ensure fairness in the reconciliation process remains an unresolved issue in the forecast reconciliation literature. This paper addresses this gap by introducing a value-oriented reconciliation approach, which we contrast with the widely studied quality-oriented reconciliation approach.
@@
To motivate our proposal and demonstrate its applicability, we focus on the well-established problem of trading wind energy in electricity markets, where a portfolio manager (PM) trades on behalf of multiple wind power producers (WPPs). Our setup closely mirrors that of Guerrero-Mestre et al., (2015). Specifically, we consider two trading floors: a forward market and a balancing market with a dual-price settlement mechanism for imbalances. As noted by Pinson et al., (2007), the trading problem for a risk-neutral, price-taker WPP is analogous to a newsvendor problem. Ultimately, it involves offering on a specific quantile of the predictive distribution for wind power generation. We assume that both the PM and WPPs are risk-neutral and act as price-takers, enabling us to model their interaction as a cooperative newsvendor game (Montrucchio et al.,, 2012). In this setup, a natural two-level hierarchy emerges between the PM and the WPPs, which implies that the PM’s offer must represent the collective output of all WPPs. As the forecasts issued by the PM and WPPs may be incoherent, they must be reconciled to ensure coherence before being used in the PM’s trading strategies. After trading in the markets, the PM allocates profits to each WPP by using pre-defined allocation rules.
@@
Our reconciliation framework is inspired by Pennings and Van Dalen, (2017), where the reconciled forecasts are expressed as the product of an aggregation matrix (defined by the hierarchy) and latent variables. The latent variables are derived from the base forecasts via learnable combination functions. Market data, including historical prices, are incorporated as context features for the combination functions, enabling the reconciliation to be adaptive to contextual information. The requirement for consensus on the reconciliation approach aligns with the group decision-making process, in which agents collaboratively select a combination function from the collection of mutually accepted functions. Particularly, we assume that WPPs exhibit individual rationality, indicating that each producer’s profit must be at least equal to their independent offering. Therefore, the forecast reconciliation process must ensure that each producer’s profit is either maintained or improved after reconciliation. To address this, we model the problem as a cooperative bargaining game (Nash,, 1953), where each agent seeks to maximize their own gain by choosing from the available combination functions. The Nash bargaining solution is derived by maximizing the product of excess profits (known as the Nash product), which measures the difference between the post-reconciliation profits and the pre-reconciliation profits.
@@
Given that the PM regularly engages in market trading, we apply the principle of fairness over time (Lodi et al.,, 2024) to determine the combination function in the reconciliation process, rather than relying on a single trade instance. Consequently, we calculate the average profit for each agent within the Nash bargaining framework. The combination function is derived through constrained optimization, with the objective of maximizing the product of average excess profits while ensuring that each agent’s average excess profit remains positive. It turns out to be a large-scale optimization problem, thus we frame it as a constrained learning problem by parameterizing the combination function (Donini et al.,, 2018). Specifically, we use the Lagrangian optimization approach outlined by Chamon and Ribeiro, (2020) for parameter estimation. The parameters of the combination function model are estimated using the stochastic gradient descent algorithm, while the dual variables are updated through the dual ascent algorithm. Finally, following Muñoz et al., (2023), we validate the proposed method with case studies in an hourly forward market coupled with a dual-price settlement balancing market, using market data from the Danish TSO and wind power data from the Wind Toolkit. In summary, this work makes two key contributions:
@@ 2 Preliminaries
In this section, we first outline the fundamentals of value-oriented forecasting and decision-making, with a specific focus on wind energy trading. We then present renewable energy offering in the forward market within the context of dual-price imbalance settlement. Finally, we describe the challenges associated with aggregated wind energy trading.
Notations: Vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), while matrices are represented by bold uppercase letters (e.g., $\mathbf{A}$). Sets are denoted using calligraphic fonts, such as $\mathcal{T}$, with the cardinality of a set expressed as $|\mathcal{T}|$. The subscript F represents the forward market (e.g., $\pi^{\texttt{F}}$), while the superscripts UP and DW denote up- and down-regulation (e.g., $\pi^{\texttt{UP}}$ and $\pi^{\texttt{DW}}$).
@@ 2.1 Value-oriented Forecasting and Decision Making
To begin, we outline the setups of wind energy offering. Let $y_{t,i}$ represent the wind energy generated by WPP $i$ at time $t$, which is a realization of the random variable $Y_{t,i}$. In a forward market, WPP $i$ must declare the energy amount $y_{t+k,i}^{\texttt{F}}$ it will generate before the market’s closure time, $t$. Here, the subscript $k$ indicates the time interval between trading and actual power delivery. In the balancing market, the WPP is required to balance the discrepancy between the actual generation, $y_{t+k,i}$, and the declared amount, $y_{t+k,i}^{\texttt{F}}$. Throughout this paper, we assume that WPPs act as price takers and are risk-neutral. This assumption is commonly observed in the electricity market literature (Morales et al.,, 2013), implying that the trading activities of a single WPP do not influence the clearing price. The prices for the forward market, upward and downward regulation are denoted by $\pi^{\texttt{F}}_{t+k}$, $\pi^{\texttt{UP}}_{t+k}$, and $\pi^{\texttt{DW}}_{t+k}$, respectively. It is established that $\pi^{\texttt{UP}}_{t+k}\geq\pi^{\texttt{F}}_{t+k}$ and $\pi^{\texttt{DW}}_{t+k}\leq\pi^{\texttt{F}}_{t+k}$, with at most one of these values differing from $\pi^{\texttt{F}}_{t+k}$.
@@
The profit of a WPP at the trading hour $t+k$ is then given by the function $\rho_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})$, which is expressed as
$\rho_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})=\pi^{\texttt{F}}_{t+k}y_{t+k,i}^{\texttt{F}}-\pi^{\texttt{UP}}_{t+k}\left[y_{t+k,i}^{\texttt{F}}-y_{t+k,i}\right]^{+}+\pi^{\texttt{DW}}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\texttt{F}}\right]^{+},$ (1)
@@
where $[x]^{+}=\max(x,0),\forall x\in\mathbb{R}$. When the power generation falls short of the offer, i.e., $y_{t+k,i}<y_{t+k,i}^{\texttt{F}}$, the WPP has to purchase up-regulation at the price of $\pi^{\texttt{UP}}_{t+k}$ to balance the discrepancy $y_{t+k,i}^{\texttt{F}}-y_{t+k,i}$. Conversely, if the power generation exceeds the offer, i.e., $y_{t+k,i}^{\texttt{F}}>y_{t+k,i}$, the WPP is required to purchase down-regulation at the price of $\pi^{\texttt{DW}}_{t+k}$. The penalties associated with overproduction and underproduction are calculated as follows: for overproduction, $\psi^{+}_{t+k}=\pi^{\texttt{F}}_{t+k}-\pi^{\texttt{DW}}_{t+k}$, and for underproduction, $\psi^{-}_{t+k}=\pi^{\texttt{UP}}_{t+k}-\pi^{\texttt{F}}_{t+k}$. It is important to observe that
$y_{t+k,i}-y_{t+k,i}^{\texttt{F}}=\left[y_{t+k,i}-y_{t+k,i}^{\texttt{F}}\right]^{+}-\left[y_{t+k,i}^{\texttt{F}}-y_{t+k,i}\right]^{+}.$
Then we recast equation (1) in the form
$\rho_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})=\pi^{\texttt{F}}_{t+k}y_{t+k,i}-\left[\psi^{+}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\texttt{F}}\right]^{+}+\psi^{-}_{t+k}\left[y_{t+k,i}^{\texttt{F}}-y_{t+k,i}\right]^{+}\right].$ (2)
@@
The first component is beyond the producer’s control, as both $\pi^{\texttt{F}}_{t+k}$ and $y_{t+k,i}$ are uncertain. The subsequent component is identified as the imbalance cost, which is a non-negative quantity represented by a function $c_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})$ as follows:
$c_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})=\psi^{+}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\texttt{F}}\right]^{+}+\psi^{-}_{t+k}\left[y_{t+k,i}^{\texttt{F}}-y_{t+k,i}\right]^{+},$ (3)
@@
Since the real generation $y_{t+k,i}$ is unknown at the time of offering $t$, WPPs must rely on forecasts to determine their offers. It is common to use the contextual information available at time $t$ such as weather and lagged values, denoted as $\mathbf{x}_{t,i}$, to predict the future value $Y_{t+k,i}$ using a specific model $f_{i}$. Denoting the point forecast as $\hat{y}_{t+k,i|t}$, forecasting can be mathematically expressed as
@@
Value-oriented forecasting: It is important to note that the optimal offer can be determined by minimizing the expected imbalance cost, provided that the distribution of $Y_{t+k,i}$ is known. It is expressed as
$y_{t+k,i}^{\texttt{F}*}=\underset{y_{t+k,i}^{\texttt{F}}\in[0,p_{i}]}{\arg\min}\ \mathbb{E}_{Y_{t+k,i}}\left[c_{t+k}(y_{t+k,i}^{\texttt{F}},y_{t+k,i})\right],$ (6)
where $p_{i}$ represents the capacity of WPP $i$. This addresses a specific instance of the newsvendor problem (Choi,, 2012), which leads to the identification of a particular quantile within the wind power generation distribution:
$y_{t+k,i}^{\texttt{F}*}=F^{-1}_{y_{t+k,i}}(\alpha_{t+k}),$ (7)
where $F_{y_{t+k,i}}$ represents the cumulative distribution function (CDF) of $Y_{t+k,i}$ and $\alpha_{t+k}$ signifies the nominal level, expressed as: $\alpha_{t+k}=\frac{\psi^{+}_{t+k}}{\psi^{+}_{t+k}+\psi^{-}_{t+k}}$. Thus, the expense associated with the imbalance is represented by $c_{t+k}(y_{t+k,i}^{\texttt{F}*},y_{t+k,i})$, and the expected cost is obtained as $\mathbb{E}_{Y_{t+k,i}}[c_{t+k}(y_{t+k,i}^{\texttt{F}*},y_{t+k,i})]$. However, the distribution of wind power generation, $F_{y_{t+k,i}}$, is unknown during the offering process. Instead, Bremnes, (2004) and Pinson et al., (2007) suggest that, for a specified nominal quantile level, it is allowed to directly estimate the optimal offers using quantile regression. Then, the parameters of the forecast model are estimated by minimizing the expected cost, expressed as
$\underset{\phi_{i}}{\min}\ \mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right].$ (8)
Specifically, Bremnes, (2004); Pinson et al., (2007) employ the long-term mean values of $\psi^{+}_{t+k}$ and $\psi^{-}_{t+k}$ during the model training process. Since the loss function aligns with the value of downstream decisions, we refer to this approach as value-oriented forecasting.

### v2

@@ Abstract
Forecast reconciliation is considered an effective method to achieve coherence (within a forecast hierarchy) and to improve forecast quality. However, the value of reconciled forecasts in downstream decision-making tasks has been mostly overlooked. In a multi-agent setup with heterogeneous loss functions, this oversight may lead to unfair outcomes, hence resulting in conflicts during the reconciliation process. To address this, we propose a value-oriented forecast reconciliation approach that focuses on the forecast value for all individual agents. Fairness is ensured through the use of a Nash bargaining framework. Specifically, we model this problem as a cooperative bargaining game, where each agent aims to optimize their own gain while contributing to the overall reconciliation process. We then present a primal-dual algorithm for parameter estimation based on empirical risk minimization. From an application perspective, we consider an aggregated wind energy trading problem, where profits are distributed using a weighted allocation rule. We demonstrate the effectiveness of our approach through several numerical experiments, showing that it consistently results in increased profits for all agents involved.
@@ 1 Introduction
The geographically distributed nature of renewable energy sources, such as wind and solar, inherently creates a hierarchical structure for power generation. It is acknowledged that forecast accuracy can be improved by leveraging the spatial-temporal dependencies inherent in geographically distributed data (tastu2013probabilistic). However, various agents at different levels of aggregation in power systems and electricity markets, for instance, the portfolio and individual power generation assets, often independently generate their own forecasts. This can lead to incoherence, i.e., the forecast directly produced at the portfolio level is not necessarily equal to the sum of the forecasts for individual assets. Given that forecasts are frequently used in decision-making, incoherence across aggregation levels may eventually lead to dilemmas among decision-makers. For that, forecast reconciliation is applied, which is a post-forecasting process that involves transforming independently generated incoherent base forecasts into coherent forecasts that satisfy the constraints imposed by the hierarchical structures (athanasopoulos2023forecast).
@@
Single-level reconciliation approaches, such as bottom-up, top-down, and middle-out approaches, have been in use for decades, dating back at least to edwards1969should and zellner1969aggregation. They require forecasts at a chosen level, which are then combined to generate forecasts for the remaining levels in the hierarchy. hyndman2011optimal introduced a combination approach that leverages forecasts across the entire hierarchy. This was later extended by van2015game, hyndman2016fast, and wickramasuriya2019optimal through employing various projection matrices (panagiotelis2021forecast). In parallel, ben2019regularized developed an end-to-end reconciliation approach, which learns the projection matrices via machine learning. Forecast reconciliation in the energy sector has been employed since van2015game, and most often brings the added benefit of enhanced forecast accuracy (zhang2018least; jeon2019probabilistic).
@@
While previous studies have substantially advanced forecast reconciliation through various methods and techniques, nearly all of them focus on leveraging hierarchical structures to improve forecast accuracy. Nonetheless, as murphy1993good suggests, in addition to evaluating the quality (accuracy) of forecasts, which is the correspondence between forecasts and observed outcomes, it is also important to assess the consistency, or the alignment between forecasts and judgments, and the value, which refers to the benefits they provide to users, in terms of profits or costs. Evidence suggests that more accurate forecasts do not always lead to higher value in downstream decision-making processes (kourentzes2013intermittent; barrow2016distributions). Although increasing emphasis is placed on developing forecast products that maximize value in decision-making, particularly within the energy industry (carriere2019integrated; stratigakos2022prescriptive; zhang2023value) and inventory management (kourentzes2020optimising; theodorou2025forecast), especially after the theoretical analysis by bertsimas2020predictive and elmachtoub2022smart, value in forecast reconciliation remains largely under-investigated, with only a few notable exceptions. For instance, sanguri2022hierarchical assessed the effectiveness of coherent forecasts in enhancing the value of port operations and empirically demonstrated that coherent forecasts improved decision-making at ports. In parallel, stratigakos2024decision proposed a method for combining multiple probabilistic forecasts using a linear opinion pool, with a particular emphasis on minimizing the expected costs associated with subsequent decision-making. However, all of these studies evaluate the value in forecast reconciliation from a centralized perspective, neglecting the fact that the agents involved may have heterogeneous loss (or cost) functions.
@@
When agents have heterogeneous cost functions, the primary challenge lies in achieving consensus on the reconciliation approach before actually reconciling their forecasts. Since the commonly used reconciliation approaches aim to optimize an overall loss function, for instance, the mean squared error function (van2015game; hyndman2016fast), some agents involved may even find that reconciled forecasts lead to an increase in their own costs, even though overall costs are reduced after reconciliation. As a result, the agents may perceive the reconciliation process as unfair, leading to disagreements on the reconciliation procedure. Indeed, studies like van2015game and sprangers2024hierarchical included weighting factors for forecasts at each level during the reconciliation process, either on an ad hoc basis or based on hierarchical levels. Nevertheless, the emphasis is primarily placed on the overall accuracy of the forecasts, thereby leaving the effects of this weighting on subsequent decision-making results unexplored. Therefore, how to account for the value of forecasts at each level and ensure fairness in the reconciliation process remains an unresolved issue in the forecast reconciliation literature. This paper addresses this gap by introducing a value-oriented reconciliation approach, which we contrast with the widely studied quality-oriented reconciliation approach.
@@
To motivate our proposal and demonstrate its applicability, we focus on the well-established problem of trading wind energy in electricity markets, where a portfolio manager (PM) trades on behalf of multiple wind power producers (WPPs). In this setup, a natural two-level hierarchy emerges between the PM and the WPPs, which implies that the PM’s offer must represent the collective output of all WPPs. As the forecasts issued by the PM and WPPs may be incoherent, they must be reconciled to ensure coherence before being used in the PM’s trading strategies. After trading in the markets, the PM allocates profits to each WPP by using pre-defined allocation rules.
@@
Our reconciliation framework is inspired by pennings2017integrated, where the reconciled forecasts are expressed as the product of a structural matrix (defined by the hierarchy) and latent variables. The latent variables are derived from the base forecasts via learnable combination functions. Market data, including historical prices, are incorporated as context features for the combination functions, enabling the reconciliation to be adaptive to contextual information. The requirement for consensus on the reconciliation approach aligns with the group decision-making process, in which agents collaboratively select a combination function from the collection of mutually accepted functions. Particularly, we assume that WPPs exhibit individual rationality, indicating that each producer’s profit must be at least equal to their independent offering. Therefore, the forecast reconciliation process must ensure that each producer’s profit is either maintained or improved after reconciliation. To address this, we model the problem as a cooperative bargaining game (nash1953two), where each agent seeks to maximize their own gain by choosing from the available combination functions. The Nash bargaining solution is derived by maximizing the product of excess profits (known as the Nash product), which measures the difference between the post-reconciliation profits and the pre-reconciliation profits.
@@
Given that the PM regularly engages in market trading, we apply the principle of fairness over time (lodi2024framework) to determine the combination function in the reconciliation process, rather than relying on a single trade instance. Consequently, we calculate the average profit for each agent within the Nash bargaining framework. The combination function is derived through constrained optimization, with the objective of maximizing the product of average excess profits while ensuring that each agent’s average excess profit remains positive. It turns out to be a large-scale optimization problem, thus we frame it as a constrained learning problem by parameterizing the combination function (donini2018empirical). Specifically, we use the Lagrangian optimization approach outlined by chamon2020probably for parameter estimation. The parameters of the combination function model are estimated using the stochastic gradient descent algorithm, while the dual variables are updated through the dual ascent algorithm. Finally, following munoz2023online, we validate the proposed method with case studies in an hourly forward market coupled with a dual-price settlement balancing market, using market data from the Danish TSO and wind power data from the Wind Toolkit. In summary, this work makes the following contributions:
@@
  3. (3)
    We prove in Corollary 3.2 that, by employing weighted proportional allocation rules, the agreement set of parameters in the Nash bargaining problem—within the context of aggregated wind energy trading—is guaranteed to be non-empty.
  4. (4)
    We propose an algorithm for parameter estimation within the Nash bargaining framework that alternates between maximizing over the dual variables and minimizing over the parameters of the reconciliation function.
@@
Notations: Vectors are denoted by bold lowercase letters (e.g., $\mathbf{x}$), while matrices are represented by bold uppercase letters (e.g., $\mathbf{A}$). Sets are denoted using calligraphic fonts, such as $\mathcal{T}$, with the cardinality of a set expressed as $|\mathcal{T}|$. The subscript $\mathrm{F}$ represents the forward market (e.g., $\pi^{\mathrm{F}}$), while the superscripts $\mathrm{UP}$ and $\mathrm{DW}$ denote up- and down-regulation (e.g., $\pi^{\mathrm{UP}}$ and $\pi^{\mathrm{DW}}$).
@@ 2 Preliminaries
Our setup closely mirrors that of guerrero2015optimal. Specifically, we consider two trading floors: a forward market and a balancing market with a dual-price settlement mechanism for imbalances. As noted by pinson2007trading, the trading problem for a risk-neutral, price-taker WPP is analogous to a newsvendor problem. Ultimately, it involves offering on a specific quantile of the predictive distribution for wind power generation. We assume that both the PM and WPPs are risk-neutral and act as price-takers, enabling us to model their interaction as a cooperative newsvendor game (montrucchio2012cooperative). In what follows, we first outline the fundamentals of value-oriented forecasting and decision-making, with a specific focus on wind energy trading. We then present renewable energy offering in the forward market within the context of dual-price imbalance settlement. Finally, we describe the challenges associated with aggregated wind energy trading.
@@ 2.1 Value-oriented Forecasting and Decision Making
To begin, we outline the setups of wind energy offering. Let $y_{t,i}$ represent the wind energy generated by WPP $i$ at time $t$, which is a realization of the random variable $Y_{t,i}$. In a forward market, WPP $i$ must declare the energy amount $y_{t+k,i}^{\mathrm{F}}$ it will generate before the market’s closure time, $t$. Here, the subscript $k$ indicates the time interval between trading and actual power delivery. In the balancing market, the WPP is required to balance the discrepancy between the actual generation, $y_{t+k,i}$, and the declared amount, $y_{t+k,i}^{\mathrm{F}}$. Throughout this paper, we assume that WPPs act as price takers and are risk-neutral. This assumption is commonly observed in the electricity market literature (morales2013integrating), implying that the trading activities of a single WPP do not influence the clearing price. The prices for the forward market, upward and downward regulation are denoted by $\pi^{\mathrm{F}}_{t+k}$, $\pi^{\mathrm{UP}}_{t+k}$, and $\pi^{\mathrm{DW}}_{t+k}$, respectively. It is established that $\pi^{\mathrm{UP}}_{t+k}\geq\pi^{\mathrm{F}}_{t+k}$ and $\pi^{\mathrm{DW}}_{t+k}\leq\pi^{\mathrm{F}}_{t+k}$, with at most one of these values differing from $\pi^{\mathrm{F}}_{t+k}$.
@@
The profit of a WPP at the trading hour $t+k$ is then given by the function $\rho_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})$, which is expressed as
$\rho_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})=\pi^{\mathrm{F}}_{t+k}y_{t+k,i}^{\mathrm{F}}-\pi^{\mathrm{UP}}_{t+k}\left[y_{t+k,i}^{\mathrm{F}}-y_{t+k,i}\right]^{+}+\pi^{\mathrm{DW}}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\mathrm{F}}\right]^{+},$ (1)
@@
where $x]^{+}=\max(x,0),\forall x\in\mathbb{R}$. When the power generation falls short of the offer, i.e., $y_{t+k,i}<y_{t+k,i}^{\mathrm{F}}$, the WPP has to purchase up-regulation at the price of $\pi^{\mathrm{UP}}_{t+k}$ to balance the discrepancy $y_{t+k,i}^{\mathrm{F}}-y_{t+k,i}$. Conversely, if the power generation exceeds the offer, i.e., $y_{t+k,i}^{\mathrm{F}}>y_{t+k,i}$, the WPP is required to purchase down-regulation at the price of $\pi^{\mathrm{DW}}_{t+k}$. The penalties associated with overproduction and underproduction are calculated as follows: for overproduction, $\psi^{+}_{t+k}=\pi^{\mathrm{F}}_{t+k}-\pi^{\mathrm{DW}}_{t+k}$, and for underproduction, $\psi^{-}_{t+k}=\pi^{\mathrm{UP}}_{t+k}-\pi^{\mathrm{F}}_{t+k}$. For mathematical convenience, we recast (1) as
$\rho_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})=\pi^{\mathrm{F}}_{t+k}y_{t+k,i}-\left[\psi^{+}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\mathrm{F}}\right]^{+}+\psi^{-}_{t+k}\left[y_{t+k,i}^{\mathrm{F}}-y_{t+k,i}\right]^{+}\right],$ (2)
given that
$y_{t+k,i}-y_{t+k,i}^{\mathrm{F}}=\left[y_{t+k,i}-y_{t+k,i}^{\mathrm{F}}\right]^{+}-\left[y_{t+k,i}^{\mathrm{F}}-y_{t+k,i}\right]^{+}.$
@@
The first term on the RHS of (2) represents the profit derived from the oracle forecast (i.e., an ideal forecast), and consequently, it is outside the producer’s influence. The subsequent term on the RHS of (2) is identified as the imbalance cost, which is a non-negative quantity represented by a function $c_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})$ as follows:
$c_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})=\psi^{+}_{t+k}\left[y_{t+k,i}-y_{t+k,i}^{\mathrm{F}}\right]^{+}+\psi^{-}_{t+k}\left[y_{t+k,i}^{\mathrm{F}}-y_{t+k,i}\right]^{+}.$ (3)
And we note that $\psi^{+}_{t+k}$ and $\psi^{-}_{t+k}$ are non-negative and bounded.
It is important to note that the optimal offer can be determined by minimizing the expected imbalance cost, provided that the distribution of $Y_{t+k,i}$ is known. It is expressed as
$y_{t+k,i}^{\mathrm{F}*}=\underset{y_{t+k,i}^{\mathrm{F}}\in[0,u_{i}]}{\arg\min}\ \mathbb{E}_{Y_{t+k,i}}\left[c_{t+k}(y_{t+k,i}^{\mathrm{F}},y_{t+k,i})\right],$ (4)
where $u_{i}$ represents the capacity of WPP $i$. This addresses a specific instance of the newsvendor problem (choi2012handbook), which leads to the identification of a particular quantile within the wind power generation distribution:
$y_{t+k,i}^{\mathrm{F}*}=F^{-1}_{y_{t+k,i}}(\alpha_{t+k}),$ (5)
where $F_{y_{t+k,i}}$ represents the cumulative distribution function (CDF) of $Y_{t+k,i}$ and $\alpha_{t+k}$ signifies the nominal level, expressed as:
$\alpha_{t+k}=\frac{\psi^{+}_{t+k}}{\psi^{+}_{t+k}+\psi^{-}_{t+k}}.$
Thus, the expense associated with the imbalance is represented by $c_{t+k}(y_{t+k,i}^{\mathrm{F}*},y_{t+k,i})$, and the expected cost is obtained as $\mathbb{E}_{Y_{t+k,i}}[c_{t+k}(y_{t+k,i}^{\mathrm{F}*},y_{t+k,i})]$.
@@
However, both the real generation $y_{t+k,i}$ and its distribution $F_{y_{t+k,i}}$ are unknown at the time of offering $t$, thus WPPs must rely on forecasts to determine their offers. It is common to use the contextual information available at time $t$ such as weather and lagged values, denoted as $\mathbf{x}_{t,i}$, to predict the future value $Y_{t+k,i}$ using a specific model $f_{i}$. Denoting the point forecast as $\hat{y}_{t+k,i|t}$, forecasting can be mathematically expressed as
@@
Value-oriented forecasting: As various studies indicate that more precise forecasts do not necessarily enhance value in subsequent decision-making (theodorou2025forecast), we can create forecasting models that directly target decision-making value. Then, the parameters of the forecast model are estimated by minimizing the expected cost, expressed as
$\underset{\phi_{i}}{\min}\ \mathbb{E}\left[c_{t+k}(\hat{y}_{t+k,i|t},y_{t+k,i})\right],$ (8)
where $c_{t+k}(\cdot)$ is defined in (3). Particularly for the wind power offering problem, bremnes2004probabilistic and pinson2007trading suggest that, for a specified nominal quantile level, it is allowed to directly estimate the optimal offers using quantile regression. Specifically, they employ the long-term mean values of $\psi^{+}_{t+k}$ and $\psi^{-}_{t+k}$ during the model training process. It is worth mentioning that the loss function is not limited to the pinball loss; it can also be a more generalized form as explored in studies by kourentzes2020optimising; stratigakos2022prescriptive; zhang2023value, which address issues related to optimization or simulation. Since the loss function aligns with the value of downstream decisions, we refer to this approach as value-oriented forecasting.